# Towards Long-Term User Welfare in Recommender Systems via Creator-Oriented Information Revelation

### v1

@@ 4. Our Approach: LoRe
In this section, to take into account creators’ bounded rationality and optimize platform’s information revelation strategy, we propose a novel method, LoRe. The overall algorithm workflow of LoRe is shown in Figure 1 and Algorithm 1. To optimize the long-term welfare of RS, we first formulate the dynamic interaction process under information asymmetry between the platform and creators as a MDP, which is described in Section 4.1. Then, we design a RL algorithm to optimize the platform’s information revelation strategy under such a dynamic environment, as presented in Section 4.2.
4.1. MDP for Platform and Creators Interactions
In our method, we use $\mathcal{A}_{p}=\{a_{i}\}_{i=0}^{|\mathcal{G}|}$ to denote the set of platform’s suggestions, where $a_{0}$ means no suggestion and $a_{i}\ (i\neq 0)$ means that platform suggests the creator to produce item with genre $g_{i}$. At the beginning of each round $t$, the platform will send suggestions selected from $\mathcal{A}_{p}$ to creators. Note that different creators may receive different suggestions. After receiving the suggestion, each creator will decide whether to follow the suggestion and choose to produce an item. Then, for each arriving user, the platform will generate $k$ items to recommend and each user can choose to click or skip these items. Then, the MDP for the interaction between platform and creators can be defined as a tuple $M=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where:
@@ 4.1. MDP for Platform and Creators Interactions
$\bm{s}^{t}=(e_{1}^{t},e_{2}^{t},\dots,e_{|\mathcal{C}|}^{t};\hat{d}_{1}^{t},\hat{d}_{2}^{t},\dots,\hat{d}_{|\mathcal{C}|}^{t})\in\mathcal{S},$
which contains the distribution of item genres and the predicted creators’ trust in round t1t-1, where $e_{i}^{t}$ denotes the genre of item created by creator $c_{i}$ and $\hat{d}_{i}^{t}$ represents the predicted trust of creator $i$ in round t1t-1. We use $\bm{e}=(e_{1},e_{2},\dots,e_{|\mathcal{C}|})$ to denote content distribution and use $\bm{\hat{d}}=(\hat{d}_{1},\hat{d}_{2},\dots,\hat{d}_{|\mathcal{C}|})$ to represent the predicted trust vector. In simple terms, we have $\bm{s}^{t}=(\bm{e}^{t};\bm{d}^{t})$.
@@
Note that we incorporate predicted creators’ trust degree into the state, which can help capture creators’ bounded rational behavior. The platform’s strategy of sending suggestion signals is defined as $\pi:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$, where $\pi(\bm{a}^{t}|\bm{s}^{t})$ is the probability of taking action $\bm{a}^{t}$ when the state at round $t$ is $\bm{s}^{t}$. Then, the expected discounted user clicks under strategy $\pi$ can be formulated as:
(1) $J(\pi)=\operatorname{\mathbb{E}}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1})\right],$
where $\tau$ is a trajectory, i.e., $\tau=(\bm{s}^{0},\bm{a}^{0},\bm{s}^{1},\bm{a}^{1},\dots)$ and $\tau\sim\pi$ is a shorthand indicating that the distribution over trajectories depends on $\pi:\bm{a}^{t}\sim\pi(\cdot|\bm{s}^{t})$, $\bm{s}^{t+1}\sim\mathcal{T}(\cdot|\bm{s}^{t},\bm{a}^{t})$. Furthermore, we respectively use $V^{\pi}(\bm{s}^{t})$, $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ and $A^{\pi}(\bm{s}^{t},\bm{a}^{t})$ to denote the state value function, state-action value function and the advantage function, where $V^{\pi}(\bm{s}^{t})$ is the value of state $\bm{s}^{t}$, $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ is the value of taking action $\bm{a}^{t}$ under state $\bm{s}^{t}$ and $A^{\pi}(\bm{s}^{t},\bm{a}^{t})=Q^{\pi}(\bm{s}^{t},\bm{a}^{t})-V^{\pi}(\bm{a}^{t})$ [^2] In practice, the advantage function $A^{\pi}(\bm{s}^{t},\bm{a}^{t})$ is calculated by the generalized advantage estimation (Schulman et al., 2015).. $V^{\pi}(s^{t})$ and $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ are common functions in MDP, and their definitions can be found in appendix B.
4.2. Reinforcement Learning Algorithm for Platform’s Strategy
In this section, we aim to optimize the long-term welfare in Eq(1) under the dynamic user-creator-platform interaction environment by learning an information-revealing strategy $\pi$. To achieve this, we design an actor-critic based RL algorithm. The overall workflow of our algorithm is shown in Algorithm 1, which contains two phases: the online inference (Line 4-13) and the offline training (Line 14-19).
@@ 4.2. Reinforcement Learning Algorithm for Platform’s Strategy
4:  // $\texttt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} Phase 1: Online Inference}}$.
@@
6:   Get item genres’ distribution $\bm{e}^{t}$ and predict trust $\hat{\bm{d}}^{t}$.
7:   Get state $\bm{s}^{t}=(\bm{e}^{t},\hat{\bm{d}}^{t})$.
8:   Sample action $\bm{a}^{t}\sim\pi_{\theta}(\bm{s}^{t})$. $\texttt{Platform sends suggestions}$.
9:   $\texttt{Creators produce items}$.
10:   The next state transitions to $\bm{s}^{t+1}$.
11:   $\texttt{Users interact with items (click or skip)}$.
12:   Get the total users’ clicks $\mathcal{R}$.
13:   Add $(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ to reply buffer $\mathcal{B}$.
14:  end for
15:  // $\texttt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} Phase 2: Offline Training}}$.
16:  for $i=1$ to $M$ do
17:   Calculate actor loss $L$ and critic loss $L^{TD}$ based on $\mathcal{B}$ according to Eq. (2) and Eq. (3).
18:   Update actor network using the gradient $\nabla_{\pi_{\theta}}L$.
19:   Update critic network using the gradient $\nabla_{V_{\phi}}L^{TD}$.
@@
4.2.1. Online Inference
At the beginning, the platform gets the current item genres’ distribution and predicts creators’ trust degree, which together constitute the state $\bm{s}^{t}$. We show how to predict creators’ trust degree in the following subsection. Then, the actor network $\pi_{\theta}$, which is parameterized by $\theta$, takes $\bm{s}^{t}$ as input and outputs the action distribution $\bm{P}^{t}=\pi_{\theta}(\bm{s}^{t})$, where the input $\bm{s}^{t}$ and output $\bm{P}^{t}$ are illustrated as follows.
Input. As shown in Figure 1, we represent the state $\bm{s}^{t}\in\mathbb{R}^{|\mathcal{C}|\times(|\mathcal{G}|+2)}$ as a one-hot matrix, where each row of the matrix is the one-hot vector of the item created by corresponding creator connected with the corresponding trust degree. The length of the one-hot vector is $|\mathcal{G}|+2$, because there are cases where the creator dose not create item.
Output. In the classic actor network, the actor usually outputs a distribution over the action space $\mathcal{A}$. However, the number of possible actions is $|\mathcal{A}|=(|\mathcal{G}|+1)^{|\mathcal{C}|}$, which is too large to handle. To address this problem, we design the actor to output the distribution over suggestion set $\mathcal{A}_{p}$ rather than the distribution over action set $\mathcal{A}$. That is, instead of outputting a joint distribution over all platform’s possible actions, we output an independent distribution over the suggested actions for each creator. Specifically, the output of actor is $\bm{P}=\pi_{\theta}(\bm{s}^{t})\in\mathbb{R}^{|\mathcal{C}|\times(\mathcal{G}|+1)}$, where each row of $\bm{P}$ is a distribution over the suggestion set for corresponding creator. Then, the suggestions sent to creators are sampled from the corresponding row of $\bm{P}$, which constitute the platform’s action $\bm{a}^{t}\in\mathbb{R}^{|\mathcal{C}|}$ as shown in Figure 1.
After sending suggestions to creators, creators produce items and the genres of these new items with creators’ trust degree constitute the next state $\bm{s}^{t+1}$. After that, platform recommends items to each user and get the total users’ clicks $\mathcal{R}$ as reward. Finally, the data $b^{t}=(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ is then added to the reply buffer $\mathcal{B}$ which is used to train the actor and critic networks in the offline training phase.
4.2.2. Offline Training
We use the collected reply buffer $\mathcal{B}$ to train the actor and critic networks. We denote the critic network by $V_{\phi}(\bm{s})$, which is parameterized by $\phi$. The critic network takes state $\bm{s}^{t}$ as input and outputs the estimated state value of $\bm{s}^{t}$, which is used to approximate the state value function $V^{\pi}(\bm{s})$ and help optimize the actor network.
@@ 4.2.2. Offline Training
(2) $L(\pi_{\theta})=-\sum_{b\in\mathcal{B}}\min\left(\pi^{\prime}(\bm{a}|\bm{s}),\text{clip}\left(\pi^{\prime}(\bm{a}|\bm{s}),1-\epsilon,1+\epsilon\right)\right)A^{\pi_{\theta}}(\bm{a},\bm{s}),$
where $\pi^{\prime}(\bm{a}|\bm{s})=\frac{\pi_{\theta}(\bm{a}|\bm{s})}{\pi_{\theta_{old}}(\bm{a}|\bm{s})}$, $\text{clip}(x,l,r)=\max(\min(x,r),l)$ and $\epsilon$ is the clip hyperparameter. $\pi_{\theta_{old}}$ is the actor network after the last update and remains unchanged when the current actor network $\pi_{\theta}$ is updated.
@@
We determine whether our algorithm has converged based on the policy loss $L(\pi_{\theta})$. Specifically, we regard the algorithm as converged when the policy loss remains nearly flat, defined as a decrease of less than 1% over 10 consecutive episodes.
4.2.3. Trust prediction
We use maximum likelihood estimation (MLE) to predict creators’ trust degrees. The training dataset $\mathcal{D}=\{(i,y)\}$ is collected from the interactions between the platform and creators, where $i$ denotes the creator’s ID, and $y\in\{0,1\}$ is a binary label indicating whether creator $i$ follows the platform’s suggestion. A neural network parameterized by $\psi$, denoted as $P_{\psi}$, is employed to model the creators’ trust. The network takes a creator’s ID as input and outputs the estimated trust degree, i.e., $\hat{d}_{i}=P_{\psi}(i)$. The parameters $\psi$ are learned by minimizing the MLE loss:
$L(P_{\psi})=-\sum_{(i,y)\in\mathcal{D}}\left[y\log(P_{\psi}(i))+(1-y)\log(1-P_{\psi}(i))\right].$
@@ 5. Experiments
In this section, we conduct extensive experiments to verify the effectiveness of our method in improving the long-term welfare of RS by addressing the following research questions:
RQ1: Can our method improve the long-term welfare of RS?
@@
RQ3: How robust our method is, and whether it can perform well under various settings?
The source code of LoRe and datasets are available  [^3] https://anonymous.4open.science/r/LoRe-E792.
@@ 5.1. Experimental Setups
Simulation setups. To better approximate real-world RS environments, we adopt the multi-stakeholder RS simulation platform proposed by Ye et al. (2025), which integrates LLM-based agents to simulate the dynamic behaviors of both users and creators. Recent studies have demonstrated that LLMs are capable of accurately mimicking user preferences and the decision-making of bounded-rational creators (Wang et al., 2024; Zhang et al., 2024; Wang et al., 2025; Ye et al., 2025). To test the robustness and generalization ability of our method under diverse environments, we also conduct additional experiments under environments where creators are simulated by non-LLM-based models (See Section 5.4). Besides, we consider the active behaviors of creators and users, as well as creator retention, following previous works (Mladenov et al., 2020; Xu et al., 2023; Ye et al., 2025): each creator $c_{i}\in\mathcal{C}$ and each user $u_{i}\in\mathcal{U}$ has probability $p_{c_{i}}$ and $p_{u_{i}}$ to be active, respectively. In each round, these probabilities determine whether a creator or user is active. Only active creators create items, and only active users receive recommendations. Moreover, a creator $c_{i}$ is assumed to leave the platform if no clicks are received in $l$ consecutive rounds of creation.
Table 2. The content diversity (CD) and average number (AN) of active creators across different methods on different datasets.
YouTube Amazon
Method CD AN CD AN
LoRe (ours) 3.259 16.250 3.462 14.710
Reweight 3.223 15.625 2.744 10.530
P-MMF 3.196 15.930 2.697 11.025
FairRec 3.182 14.245 2.646 9.880
Most History Click 3.197 14.730 2.805 11.610
Most Click 3.186 13.840 2.847 11.775
w/o signal 3.177 11.750 2.546 9.665
@@
Baselines. We compared our method with several baselines, including: w/o Signal: revealing no information to the creator at all, except for their own creation history; and some naive information revelation strategies: Most History Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres the creator has ever created; Most Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres. In addition, we compared our method with approaches that focus on recommendation algorithm design, such as P-MMF (Xu et al., 2023) and FairRec (Guo et al., 2023). We also included a baseline that influences creators by adjusting their rewards, namely Reweight (Yao et al., 2024).
@@
Metrics. Following previous works (Ye et al., 2025; Yao et al., 2024), we use the cumulative user click number to measure the long-term welfare. We adopt RecAgent (Wang et al., 2025) as the user simulator, where users can choose to click or skip when presented with an item. The number of clicks by all users in each round is recorded as the user welfare for that round.
@@
Figure 3. The number of genres created by each creator.

### v2

@@ 4. Our Approach: LoRe
In this section, to address the challenge by applying the framework of cheap talk into RS and optimizing the platform’s suggestion strategy, we propose a novel method, LoRe. The overall algorithm workflow of LoRe is shown in Figure 1 and Algorithm 1. We first formulate the dynamic interaction process between the platform and creators under information asymmetry as a Markov decision process (MDP), which is described in Section 4.1. We then design a reinforcement learning algorithm to optimize the platform’s suggestion strategy in this dynamic environment, as presented in Section 4.2.
4.1. Interactions Between Platform and Creators
We now formulate the dynamic interaction between the platform and creators in each round as follows. At the beginning of round $t$, the platform observes the state $s^{t}\in\mathcal{S}$, which represents the distribution of item genres. The platform then sends each creator a suggestion from suggestions set $\mathcal{A}_{p}$. Upon receiving a suggestion, a creator can choose to produce an item with any possible genre $g_{i}\in\mathcal{G}$. We do not impose any assumption of creator’s creation behavior, which indicates that any creator’s behavior model can be embedded in our method. After content creation and recommendation, creator $c_{i}$ obtains the number of clicks of his item as utility, denoted by $\mu_{c_{i}}^{t}$. Then, the cumulative utility of creator $c_{i}$ over $T$ rounds is given by:
$\mu_{c_{i}}=\sum_{t=1}^{T}\mu_{c_{i}}^{t}.$
The platform’s objective is to maximize the long-term user welfare, measured by the total number of user clicks, through sending appropriate suggestions. With a slight abuse of notation, we use $\bm{a}^{t}=(a_{1}^{t},a_{2}^{t},\dots,a_{|\mathcal{C}|}^{t})$ to denote the platform’s suggestion strategy in round $t$, where $a_{i}^{t}\in\mathcal{A}_{p}$ is the suggestion sent to creator $c_{i}$. Then, the platform’s problem can be formulated as follows:
$\max_{(\bm{a}^{1},\bm{a}^{2},\dots,\bm{a}^{T})}\sum_{c\in\mathcal{C}}\mu_{c}.$
This interaction can be abstracted as follows. Given a state, the platform sends suggestions to creators, after which the system transitions to a new state as a result of creators’ creation behavior, and the platform receives a reward equal to the total number of user clicks. Moreover, because the state is defined by the distribution of item genres, the state transition can be reasonably approximated as depending only on the current state and the platform’s actions. Consequently, the dynamic interaction between the platform and creators can be naturally modeled as a Markov decision process. We formulate this MDP between the interaction between platform and creators as a tuple $M=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where:
@@ 4.1. Interactions Between Platform and Creators
$\bm{s}^{t}=(e_{1}^{t},e_{2}^{t},\dots,e_{|\mathcal{C}|}^{t})\in\mathcal{S},$
which is the distribution of item genres in round t1t-1, where $e_{i}^{t}$ denotes the genre of item created by creator $c_{i}$ in round t1t-1.
@@
The platform’s suggestion strategy is redefined as:
$\pi_{\theta}:\mathcal{S}\times\mathcal{A}\mapsto[0,1],$
where $\pi_{\theta}(\bm{a}^{t}|\bm{s}^{t})$ is the probability of taking action $\bm{a}^{t}$ when the state at round $t$ is $\bm{s}^{t}$. Then, the expected discounted user clicks under strategy $\pi_{\theta}$ can be formulated as:
(1) $J(\pi_{\theta})=\operatorname{\mathbb{E}}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1})\right],$
where $\tau$ is a trajectory, i.e., $\tau=(\bm{s}^{0},\bm{a}^{0},\bm{s}^{1},\bm{a}^{1},\dots)$ and $\tau\sim\pi_{\theta}$ is a shorthand indicating that the distribution over trajectories depends on $\pi_{\theta}:\bm{a}^{t}\sim\pi_{\theta}(\cdot|\bm{s}^{t})$, $\bm{s}^{t+1}\sim\mathcal{T}(\cdot|\bm{s}^{t},\bm{a}^{t})$. Furthermore, we respectively use $V^{\pi_{\theta}}(\bm{s}^{t})$, $Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ and $A^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ to denote the state value function, state-action value function and the advantage function, where $V^{\pi_{\theta}}(\bm{s}^{t})$ is the value of state $\bm{s}^{t}$, $Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ is the value of taking action $\bm{a}^{t}$ under state $\bm{s}^{t}$ and $A^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})=Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})-V^{\pi_{\theta}}(\bm{a}^{t})$ [^1] In practice, the advantage function $A^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ is calculated by the generalized advantage estimation (Schulman et al., 2015).. $V^{\pi_{\theta}}(s^{t})$ and $Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})$ are common functions in MDP, and their definitions are as follows:
$V^{\pi_{\theta}}(\bm{s}^{t})=\operatorname{\mathbb{E}}_{(\bm{a}^{t},\bm{s}^{t+1},\dots)\sim\pi_{\theta}}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right],$
$Q^{\pi_{\theta}}(\bm{s}^{t},\bm{a}^{t})=\operatorname{\mathbb{E}}_{(\bm{s}^{t+1},\bm{a}^{t+1},\dots)\sim\pi_{\theta}}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right].$
4.2. Reinforcement Learning Algorithm for Platform’s Suggestion Strategy
In this section, we aim to optimize the long-term welfare in Eq(1) under the dynamic user-creator-platform interaction environment by learning an suggestion strategy $\pi_{\theta}$. To achieve this, we design an actor-critic based RL algorithm. The overall workflow of our algorithm is shown in Algorithm 1, which contains two phases: the inference phase (Line 4-13) and the training phase (Line 14-19).
@@ 4.2. Reinforcement Learning Algorithm for Platform’s Suggestion Strategy
4:  // $\texttt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} Inference Phase}}$.
@@
6:   Get state $\bm{s}^{t}$.
7:   Sample action $\bm{a}^{t}\sim\pi_{\theta}(\bm{s}^{t})$. $\texttt{Platform sends suggestions}$.
8:   $\texttt{Creators produce items}$.
9:   The next state transitions to $\bm{s}^{t+1}$.
10:   $\texttt{Users interact with items (click or skip)}$.
11:   Get the total users’ clicks $\mathcal{R}$.
12:   Add $(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ to reply buffer $\mathcal{B}$.
13:  end for
14:  // $\texttt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} Training Phase}}$.
15:  for $i=1$ to $M$ do
16:   Calculate actor loss $L$ and critic loss $L^{TD}$ based on $\mathcal{B}$ according to Eq. (2) and Eq. (3).
17:   Update actor network using the gradient $\nabla_{\pi_{\theta}}L$.
18:   Update critic network using the gradient $\nabla_{V_{\phi}}L^{TD}$.
19:  end for
@@
4.2.1. Inference Phase
At the beginning, the platform gets the current item genres’ distribution, which constitute the state $\bm{s}^{t}$. Then, the actor network $\pi_{\theta}$, which is parameterized by $\theta$, takes $\bm{s}^{t}$ as input and outputs the action distribution:
$\bm{P}^{t}=\pi_{\theta}(\bm{s}^{t}).$
where the input $\bm{s}^{t}$ and output $\bm{P}^{t}$ are illustrated as follows.
Input. As shown in Figure 1, we represent the state $\bm{s}^{t}\in\mathbb{R}^{|\mathcal{C}|\times(|\mathcal{G}|+1)}$ as a one-hot matrix, where each row of the matrix is the one-hot vector of the item created by corresponding creator. The length of the one-hot vector is $|\mathcal{G}|+1$, because there are cases where the creator dose not create item.
Output. In the classic actor network, the actor usually outputs a distribution over the action space $\mathcal{A}$. However, the number of possible actions is $|\mathcal{A}|=(|\mathcal{G}|+1)^{|\mathcal{C}|}$, which is too large to handle. To address this problem, we design the actor to output the distribution over suggestion set $\mathcal{A}_{p}$ rather than the distribution over action set $\mathcal{A}$. That is, instead of outputting a joint distribution over all platform’s possible actions, we output an independent distribution over the suggested actions for each creator. Specifically, the output of actor is $\bm{P}=\pi_{\theta}(\bm{s}^{t})\in\mathbb{R}^{|\mathcal{C}|\times(\mathcal{G}|+1)}$, where each row of $\bm{P}$ is a distribution over the suggestion set for corresponding creator. The suggestions sent to creators are sampled from the corresponding row of $\bm{P}$, which constitute the platform’s action $\bm{a}^{t}\in\mathbb{R}^{|\mathcal{C}|}$.
After sending suggestions to creators, creators produce items and the genres of these new items constitute the next state $\bm{s}^{t+1}$. After that, platform recommends items to each user and get the total users’ clicks $\mathcal{R}$ as reward. Finally, the data $b^{t}=(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ is then added to the reply buffer $\mathcal{B}$ which is used to train the actor and critic networks in the training phase.
4.2.2. Training Phase
We use the collected reply buffer $\mathcal{B}$ to train the actor and critic networks. We denote the critic network by $V_{\phi}(\bm{s})$, which is parameterized by $\phi$. The critic network takes state $\bm{s}^{t}$ as input and outputs the estimated state value of $\bm{s}^{t}$, which is used to approximate the state value function $V^{\pi_{\theta}}(\bm{s})$ and help optimize the actor network.
@@ 4.2.2. Training Phase
(2) $L(\pi_{\theta})=-\sum_{b\in\mathcal{B}}\min\left(\pi_{\theta}^{\prime}(\bm{a}|\bm{s}),\text{clip}\left(\pi_{\theta}^{\prime}(\bm{a}|\bm{s}),1-\epsilon,1+\epsilon\right)\right)A^{\pi_{\theta}}(\bm{a},\bm{s}),$
where $\pi_{\theta}^{\prime}(\bm{a}|\bm{s})=\frac{\pi_{\theta}(\bm{a}|\bm{s})}{\pi_{\theta_{old}}(\bm{a}|\bm{s})}$, $\text{clip}(x,l,r)=\max(\min(x,r),l)$ and $\epsilon$ is the clip hyperparameter. $\pi_{\theta_{old}}$ is the actor network after the last update and remains unchanged when the current actor network $\pi_{\theta}$ is updated.
@@
We assess convergence based on the policy loss $L(\pi_{\theta})$. Specifically, the algorithm is considered converged when the policy loss remains approximately stable, operationalized as a decrease of less than 1% over 10 consecutive episodes.
@@ 5. Experiments
In this section, we conduct extensive experiments to evaluate the effectiveness of our method in improving the long-term welfare of RS. Specifically, we aim to address the following research questions::
RQ1: Can our method improve the long-term welfare of RS through information revelation?
@@
RQ3: How robust is our method, and how does it perform under various settings?
The source code of LoRe and the datasets used in our experiments are publicly available at https://anonymous.4open.science/r/LoRe-E5F0.
@@ 5.1. Experimental Setups
Simulation setups. To better simulate the real-world RS environments, we adopt the multi-stakeholder RS simulation platform proposed by Ye et al. (2025), which integrates LLM-based agents to simulate the dynamic behaviors of both users and creators. Recent studies have demonstrated that LLMs are capable of accurately mimicking user preferences and the decision-making of bounded-rational creators (Wang et al., 2024; Zhang et al., 2024; Wang et al., 2025; Ye et al., 2025). To evaluate the robustness and generalization ability of our method under diverse environments, we additionally conduct experiments under environments where creators are simulated using non-LLM-based models (See Section 5.4). Following previous works (Mladenov et al., 2020; Xu et al., 2023; Ye et al., 2025), we also consider the active behaviors of creators and users, as well as creator retention. Specifically, each creator $c_{i}\in\mathcal{C}$ and each user $u_{i}\in\mathcal{U}$ has probability $p_{c_{i}}$ and $p_{u_{i}}$ to be active, respectively. In each round, these probabilities determine whether a creator or user is active. Only active creators create items, and only active users receive recommendations. Moreover, a creator $c_{i}$ is assumed to leave the platform if no clicks are received in $l$ consecutive rounds of creation.
Table 2. Comparisons of average user clicks, content diversity and active creators of our methods and baselines on YouTube and Amazon datasets. MHC represents the most history click and MC denotes the most click. We categorize baselines into three types: naive suggestion strategy, re-ranking methods, and reward adjust method.
User Clicks Content Diversity Active Creators
Methods YouTube Amazon YouTube Amazon YouTube Amazon
w/o Signal $101.211\pm 3.941$ $29.55\pm 0.896$ $3.177\pm 0.006$ $2.546\pm 0.021$ $11.750\pm 0.153$ $9.665\pm 0.074$
Naive MHC $105.831\pm 3.058$ $57.913\pm 1.828$ $3.197\pm 0.012$ $2.805\pm 0.006$ $4.730\pm 0.214 11$$1.610\pm 0.175$
[…]
Re-Ranking P-MMF $101.599\pm 1.263$ $47.794\pm 1.114$ $3.196\pm 0.009$ $2.697\pm 0.011$ $5.230\pm 0.120 11$$1.025\pm 0.097$
FairRec $97.171\pm 3.396$ $38.860\pm 0.747$ $3.182\pm 0.011$ $2.646\pm 0.009$ $14.245\pm 0.153$ $9.880\pm 0.068$
Reward Adjust Reweight $\underline{110.891}\pm 2.505$ $47.412\pm 1.290$ $\underline{3.223}\pm 0.004$ $2.744\pm 0.013$ $\underline{15.625}\pm 0.127$ $10.530\pm 0.082$
Ours LoRe $\bm{127.113}\pm 2.589$ $\bm{81.36}\pm 0.981$ $\bm{3.259}\pm 0.012$ $\bm{3.462}\pm 0.007$ $\bm{16.250}\pm 0.148$ $\bm{14.710}\pm 0.164$
@@
Baselines. We compared our method with several baselines, including the following: w/o Signal: revealing no information to the creator at all, except for their own creation history. Some naive suggestion strategies are as follows. Most History Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres the creator has ever created; Most Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres. In addition to these naive suggestion strategies, we compared our method with approaches that focus on recommendation algorithm design, such as P-MMF (Xu et al., 2023) and FairRec (Guo et al., 2023). We also included a baseline that influences creators by adjusting their rewards rather than information revelation, namely Reweight (Yao et al., 2024).
@@
Metrics. Following previous works (Ye et al., 2025; Yao et al., 2024), we measure the long-term welfare of RS using the cumulative number of user clicks. We adopt RecAgent (Wang et al., 2025) as the user simulator, in which users decide whether to click on or skip a presented item. The total number of clicks across all users in each round is recorded as the user welfare for that round.
@@
Figure 3. Number of genres created per creator. The blue box denotes the mean number of genres created by each creator under LoRe, while the red box denotes the mean under the w/o Signal baseline.